# Skeleton-in-Context: Unified Skeleton Sequence Modeling with In-Context Learning

### v1

@@ Abstract
In-context learning provides a new perspective for multi-task modeling for vision and NLP. Under this setting, the model can perceive tasks from prompts and accomplish them without any extra task-specific head predictions or model finetuning. However, Skeleton sequence modeling via in-context learning remains unexplored. Directly applying existing in-context models from other areas onto skeleton sequences fails due to the inter-frame and cross-task pose similarity that makes it outstandingly hard to perceive the task correctly from a subtle context. To address this challenge, we propose Skeleton-in-Context (SiC), an effective framework for in-context skeleton sequence modeling. Our SiC is able to handle multiple skeleton-based tasks simultaneously after a single training process and accomplish each task from context according to the given prompt. It can further generalize to new, unseen tasks according to customized prompts. To facilitate context perception, we additionally propose a task-unified prompt, which adaptively learns tasks of different natures, such as partial joint-level generation, sequence-level prediction, or 2D-to-3D motion prediction. We conduct extensive experiments to evaluate the effectiveness of our SiC on multiple tasks, including motion prediction, pose estimation, joint completion, and future pose estimation. We also evaluate its generalization capability on unseen tasks such as motion-in-between. These experiments show that our model achieves state-of-the-art multi-task performance and even outperforms single-task methods on certain tasks.
@@ 1 Introduction
The idea of multi-task models has attracted much research interest over the years in a wide range of areas [6, 59, 15, 16], such as computer vision and natural language processing (NLP). To build such models, a prominent line of works focus on deploying task-specific heads on a shared backbone and learning them through pretraining and finetuning. Models with such task-specific architectures are not learned in an end-to-end manner and are hard to generalize to new, unseen tasks and data. On skeleton-based human-centric tasks, most existing methods rely on task-specific designs and/or non-end-to-end frameworks to accomplish multiple tasks. Several works [9, 36] learn unified human motion representations. MotionBERT [76] adopts two-stage training, including pre-training on the 2D-3D lifting task and fine-tuning on each task. UniHCP [12] connects a well-designed shared backbone network to specific task heads according to different tasks. UPS [16] unifies action labels and motion sequences into text-based language sequences and fine-tunes them on each task for better performance. These models can have a limited scope of tasks they are capable of handling and lack of flexibility, which limits their capability to generalize to unseen tasks and data. To overcome these challenges, in-context learning represents a new trend to build generalist models, which is originally proposed in NLP [6, 48, 47] and then introduced into other areas such as images and point clouds. Some generalist models [5, 59, 60, 71] utilize the Masked Image Modeling (MIM) framework to exhibit multi-task capability via in-context learning in image processing, as shown in Fig. 1(a). Similarly, in Fig. 1(b), Point-In-Context [15] proposes a key joint sampling module and introduces in-context learning into point cloud processing. Specifically, in-context models learn patterns of provided input-output pairs, called in-context examples or prompts, and then perform the desired task according to the prompt. An in-context model can accomplish diverse tasks end-to-end without any task-specific heads or fine-tuning. It also has excellent generalization capability. Directly adapting existing in-context models from other areas onto skeleton sequences fails because they largely focus on static objects (images or point clouds) and thus cannot capture spatial-temporal dependencies in skeleton sequences. To the best of our knowledge, in-context learning remains unexplored in 3D skeleton sequence modeling.
@@ 2 Related Work
Skeleton-Based Motion Analysis. Many of the skeleton-based human-centric tasks have shared similar development patterns in terms of the methods used to address them. On human motion prediction, early successes are achieved with RNNs and CNNs [41, 27]. Nowadays, the task is dominated by Graph Convolutional Networks (GCNs). LTD [39] considers each joint coordinate as a node in a graph on which to employ learnable spatial graph convolutions. Follow-up works [40, 13] focus on how to assist graph convolutions with additional human structural information or patterns. Recent works [28, 37] combine spatial and temporal graph convolutions to learn spatial-temporal dependencies. Different from motion prediction, the goal of motion synthesis is to complete the human motion sequence. To complete the missing parts of human motion, many works adopt convolutional models [67, 25, 21]or generative adversarial networks [75, 7, 19, 22]. one category of 3D pose estimation [17, 32, 34] is to lift the estimated 2D pose to 3D, which often employs spatial-temporal convolution [42, 46, 10, 8, 11, 58, 66] or transformer-based [70, 73, 51, 72, 74] methods. However, these approaches focus on designing customized architectures specific to a task, which are adaptive to a limited range of applications. To this end, we propose to unify the input and output formats of multiple tasks and utilize a single model to process multiple tasks simultaneously via in-context learning.
@@
In-Context Learning. The emergence of in-context learning provides a new paradigm for achieving multi-tasking, which originated from Natural Language Processing (NLP) [48, 47, 14, 49, 1]. It endows models with the capability to handle multiple tasks and even unseen tasks by incorporating domain-specific input-target pairs, known as prompts, into the query sample [6]. Following NLP, in-context learning is extended to computer vision in some work, such as 2D images [59, 60, 53, 71] and 3D point clouds [63, 3, 15]. Specifically, Visual Prompt [5], Painter [59], and Point-In-Context [15] are pioneers of in-context methods in computer vision. They verify the effectiveness and flexibility of in-context learning in vision-based tasks, and most of them adopt pure-transformer-based approaches [20, 4, 29, 61]. Meanwhile, several approaches deeply explore Prompt Engineering [53, 71, 3] and awe-inspiring generalization [60, 15, 63] in in-context learning. Recently, LVM [2] unifies most vision tasks and data via joint co-training on one visual in-context model. Compared to the works focused on static objects (images and point clouds), our work mainly delves into modeling in-context dynamic skeleton sequences.
@@
Unified Skeleton-Centric Models. A generalist model that can excel in handling multiple tasks is a profound step toward general artificial intelligence [59, 47]. Previous works [24, 26, 52, 55, 31, 65, 35, 68] give the effort to explore co-training several human-centric tasks. Meanwhile, some works [30, 44] show great progress in simultaneously predicting parsing and pose. Recent research [9, 36] demonstrates the effectiveness of modeling unified human motion sequences. In particular, UniHCP [12] designs specific task heads according to different 2D human-centric tasks. Motion-BERT [76] is pre-trained in the 2D-3D lifting task and fine-tuned on each task. UPS [16] unifies heterogeneous human behavior understanding tasks via language sequences and cycle training for each task in order. Nevertheless, these works are limited by the dilemma of complex training stages or task heads, which introduce extra costs. We integrate the output format of different tasks in skeleton sequence format and train the model for just one time.
@@ 3.1 Formulation
where $k$ represents the task index while $i$ and $j$ indicate different example indexes. Thus, we can determine what task will be performed on the query sample by selecting the task-specific prompt from the corresponding training set.
@@
Figure 3: Overall framework of our Skeleton-in-Context. Specifically, we establish a skeleton bank by integrating training sets under different tasks, which contain a large amount of input-target pairs performing different tasks. Next, we randomly select a sample pair as the task-guided prompt (TGP) and a query input from the skeleton bank, undergo encoding, and concatenating, respectively, and then input them into the transformer in parallel. In particular, during this process, the query input and task-unified prompt (TUP) are combined to form a new query. After iterating $n_{1}$ times, the TGP and query are aggregated through $\mathcal{A}\left(\cdot\right)$ and then input into the transformer for $n_{2}$ iterations. Lastly, the second half of the model output is used as our prediction.
@@ 3.2 Task Definitions
To measure multitasking capability, we establish a new in-context benchmark from common skeleton-based datasets including H3.6M [23], AMASS [38], and 3DPW [57]. In this benchmark, we select three common tasks: motion prediction, pose estimation, and joint completion, and a new task: future pose estimation. Consequently, we conduct extensive experiments on a large and diverse dataset collection, which consists of 176K skeleton sequence samples.
Motion Prediction. The goal of motion prediction is to predict future motion sequences based on historical motion sequences. There is continuity between the input and output sequences, which constitute a complete motion. Thus, the input $I$ and target $T$ are complete skeleton sequences with shapes of $(F,J,3)$.
Pose Estimation. One category of 3D pose estimation [17, 32, 34] is lifting the estimated 2D pose that contains noise to 3D pose. The model takes the detected 2D skeleton sequence $I\in\mathbb{R}^{F\times J\times 2}$ as input, and estimates the correct 3D skeleton sequence $T\in\mathbb{R}^{F\times J\times 3}$. Thus, the goal of the pose estimation comprises denoising and prediction. Note that we expanded the input sequence to 3 dimensions by setting the third dimension to 00, in order to unify the data format with other tasks. Therefore, the input skeleton sequence becomes the same shape $(F,J,3)$ as the target sequence.
Future Pose Estimation. Based on motion prediction and pose estimation, we propose a new task called future pose estimation. Generally, 2D images are easier to obtain than 3D data, so once we can predict future 3D skeleton sequences based on historical 2D skeleton sequences, it will be very meaningful and bring the human perception community one step forward. Unlike traditional motion prediction, future pose estimation takes less information to predict future pose. We keep the format of input-target uniform as the above tasks. Given a historical 2D skeleton sequence $I\in\mathbb{R}^{F\times J\times 3}$ (one dimension with 00), future pose estimation expects the model to output its future 3D skeleton sequence $T\in\mathbb{R}^{F\times J\times 3}$. To our knowledge, future pose estimation has never been proposed by any works.
@@ 4 Skeleton-in-Context
${}^{\dagger}$ [42] WACV’23 19.0 30.0 35.3 46.8 52.4 36.7 74.6 60.0 59.0 67.4 63.2 64.4 68.4 66.4
${}^{\dagger}$ [64] CVPR’23 14.0 25.3 30.2 40.5 46.6 31.3 163.3 110.1 38.9 56.0 47.5 88.1 89.1 88.6
${}^{\dagger}$ [54] CVPR’23 17.4 28.0 31.5 43.0 40.4 33.9 55.6 45.6 33.6 50.6 42.1 62.2 71.7 67.0
${}^{\dagger}$ [69] ICCV’23 35.3 39.8 42.3 49.2 56.1 44.5 78.6 62.9 74.1 81.9 78.0 60.6 68.1 64.4
${}^{\ddagger}$  [76] ICCV’23 14.3 22.8 26.7 36.4 41.7 28.4 55.8 45.7 31.3 46.7 39.0 75.5 96.8 86.2
@@
4.1 Skeleton Sequence Prompt
Task-Guided Prompt. In order to endow the model with the ability to handle multiple tasks simultaneously, a prompt is needed to instruct the model on what task should be conducted on the Query sample. This is exactly the meaning of in-context learning: learning the potential pattern hidden in the Prompt from the contextual input. Here, the potential pattern contained between the input and output of the Prompt is the task pattern. As mentioned in Sec. 3.1, we propose the Task-Guided Prompt (TGP) $G^{k}_{i}=\left\{I^{k}_{i},T^{k}_{i}\right\}$, which comprises of an example skeleton pair performing a specific task $k$. In this way, our proposed model is able to perform a variety of tasks according to the provided prompts that could be seen or even unseen in the training set. It is worth that our model can perform unseen tasks and transfer to datasets that are different styles from the training dataset. It will be elaborated in Sec. 5.2.
Avoiding Over-Fitting. Previous works have great progress in unifying different tasks via the MIM-style framework [5, 59, 15]. Compared with static objects in 2D and 3D, such as images and point clouds, the obstacle to in-context learning of skeleton sequences lies in the wondrous similarity of each frame, so directly using the random masking strategy [59, 5, 15] will cause the model to fall into over-fitting, leading it to only learn simple interpolation frames rather than capturing meaningful motion sequence representations during the training stage. As shown in Fig. 2, when we directly utilize the MPM-style [20, 45] framework to train our model, although it can perfectly reconstruct the masked frame during training, the experimental results reveal that it outputs a devastating result during inference. It is obvious that the output skeleton sequence gradually shrinks as the length increases because the next frame of the current frame is 0, and interpolating with 0 will lead to shrinking. Such a phenomenon underscores the challenges of learning contextual information about dynamic skeleton sequences.
Task-Unified Prompt. With the Task-Guided Prompt that exhibits unique capabilities specific to in-domain or out-of-domain tasks, mitigating the gap between different tasks brings a new challenge. Meanwhile, as discussed above, a straight extension of previous works will lead to a simple model that can only perform the interpolation task based on the given skeleton sequence. To tackle these challenges, we introduce the Task-Unified Prompt (TUP), a well-designed and task-agnostic module, which can adaptively learn to incorporate multiple tasks into the framework of in-context learning, i.e., a one-off end-to-end framework for all tasks without having to deploy task-specific heads or fine-tuning. We introduce two versions of TUP:
@@ 4.1 Skeleton Sequence Prompt
The SiC based on Static-/Dynamic-TUP is referred to as Staitc-/Dynamic-SiC respectively. The proposed TUP facilitates the modeling of distribution differences among tasks and datasets, so the model focuses more on learning unified representations shared across tasks.
@@ 4.2 Model Architecture
As shown in Fig. 3, given a prompt pair $G^{k}_{i}=\left\{I^{k}_{i},T^{k}_{i}\right\}$ and a query sample $Q^{k}_{j}$, we first map them to the C-dimensional feature space, and then splice $I$ and $T$, $Q$ and $U$ respectively. The former represents our proposed Task-Guided Prompt, and the latter is the objective sequence of our focused analysis. The process can be formulated as:

### v2

@@ Abstract
In-context learning provides a new perspective for multi-task modeling for vision and NLP. Under this setting, the model can perceive tasks from prompts and accomplish them without any extra task-specific head predictions or model fine-tuning. However, skeleton sequence modeling via in-context learning remains unexplored. Directly applying existing in-context models from other areas onto skeleton sequences fails due to the similarity between inter-frame and cross-task poses, which makes it exceptionally hard to perceive the task correctly from a subtle context. To address this challenge, we propose Skeleton-in-Context (SiC), an effective framework for in-context skeleton sequence modeling. Our SiC is able to handle multiple skeleton-based tasks simultaneously after a single training process and accomplish each task from context according to the given prompt. It can further generalize to new, unseen tasks according to customized prompts. To facilitate context perception, we additionally propose a task-unified prompt, which adaptively learns tasks of different natures, such as partial joint-level generation, sequence-level prediction, or 2D-to-3D motion prediction. We conduct extensive experiments to evaluate the effectiveness of our SiC on multiple tasks, including motion prediction, pose estimation, joint completion, and future pose estimation. We also evaluate its generalization capability on unseen tasks such as motion-in-between. These experiments show that our model achieves state-of-the-art multi-task performance and even outperforms single-task methods on certain tasks.
@@ 1 Introduction
The idea of multi-task models has attracted much research interest over the years in a wide range of areas [6, 59, 16, 17], such as computer vision and natural language processing (NLP). To build such models, most previous works focus on deploying task-specific heads on a shared backbone and learning them through pretraining and fine-tuning. Models with such task-specific architectures are not learned in an end-to-end manner and are hard to generalize to new, unseen tasks and data. Most existing methods for skeleton-based human-centric tasks rely on task-specific designs and/or non-end-to-end frameworks to accomplish multiple tasks. Several works [9, 36] learn unified human motion representations. In particular, MotionBERT [79] adopts two-stage training, including pre-training on the 2D-3D lifting task and fine-tuning on each task. UniHCP [12] connects a well-designed shared backbone network to specific task heads according to different tasks. Meanwhile, UPS [17] unifies action labels and motion sequences into text-based language sequences and fine-tunes them on each task for better performance. These models can have a limited scope of tasks they are capable of handling and lack flexibility. This limits their capability to generalize to unseen tasks and data. To overcome these challenges, in-context learning represents a new trend to build generalist models, which was originally proposed in NLP [6, 47, 46] and then introduced into other areas such as images and point clouds. Some generalist models [5, 59, 60, 74] utilize the Masked Image Modeling (MIM) framework to exhibit multi-task capability via in-context learning in image processing, as shown in Fig. 1(a). Similarly, in Fig. 1(b), Point-In-Context [16] proposes a key joint sampling module and introduces in-context learning into point cloud processing. Specifically, in-context models learn patterns of provided input-output pairs, called in-context examples or prompts, and then perform the desired task according to the prompt. An in-context model can accomplish diverse tasks end-to-end without any task-specific heads or fine-tuning. It also has excellent generalization capability. Directly adapting existing in-context models from other areas onto skeleton sequences fails because they largely focus on static objects (images or point clouds) and thus cannot capture spatial-temporal dependencies in skeleton sequences. To the best of our knowledge, in-context learning remains unexplored in 3D skeleton sequence modeling.
@@ 2 Related Work
Skeleton-Based Motion Analysis. Many of the skeleton-based human-centric tasks have shared similar development patterns in terms of the methods used to address them. On human motion prediction, early successes are achieved with RNNs and CNNs [40, 28]. Nowadays, the task is dominated by Graph Convolutional Networks (GCNs). LTD [38] considers each joint coordinate as a node in a graph on which to employ learnable spatial graph convolutions. Follow-up works [39, 13, 61, 62] focus on designing better graph convolution networks. Different from motion prediction, the goal of motion synthesis is to complete the human motion sequence. To complete the missing parts of human motion, many works adopt convolutional models [70, 26, 22]or generative adversarial networks [78, 7, 20, 23]. One category of 3D pose estimation [18, 32, 34] is to lift the estimated 2D pose to 3D, which often employs spatial-temporal convolution [41, 45, 10, 8, 11, 58, 69] or transformer-based [73, 76, 50, 75, 77, 14] methods. However, these approaches focus on designing customized architectures specific to a task, which are limited in range of applications. To this end, we propose to unify the input and output formats of multiple tasks and utilize a single model to process multiple tasks simultaneously via in-context learning.
@@
In-Context Learning. The emergence of in-context learning provides a new paradigm for achieving multi-tasking, which originated from Natural Language Processing (NLP) [47, 46, 15, 48, 1]. It endows models with the capability to handle multiple tasks and even unseen tasks by incorporating domain-specific input-target pairs, known as prompts, into the query sample [6]. Following NLP, in-context learning is extended to computer vision in some work, such as 2D images [59, 60, 52, 74, 63, 57] and 3D point clouds [66, 3, 16]. Specifically, Visual Prompt [5], Painter [59], and Point-In-Context [16] are pioneers of in-context methods in computer vision. They verify the effectiveness and flexibility of in-context learning in vision-based tasks, and most of them adopt pure-transformer-based approaches [21, 4, 29, 64]. Meanwhile, several approaches deeply explore Prompt Engineering [52, 74, 3] and awe-inspiring generalization [60, 16, 66] in in-context learning. Recently, LVM [2] unifies most vision tasks and data via joint co-training on one visual in-context model. Compared to the works focused on static objects (images and point clouds), our work mainly delves into modeling in-context dynamic skeleton sequences.
@@
Unified Skeleton-Centric Models. A generalist model that can excel in handling multiple tasks is a profound step toward general artificial intelligence [59, 46]. Previous works [25, 27, 51, 54, 31, 68, 35, 71] give the effort to explore co-training several human-centric tasks. Meanwhile, some works [30, 43] show great progress in simultaneously predicting parsing and pose. Recent research [9, 36] demonstrates the effectiveness of modeling unified human motion sequences. In particular, UniHCP [12] designs specific task heads according to different 2D human-centric tasks. Motion-BERT [79] is pre-trained in the 2D-3D lifting task and fine-tuned on each task. UPS [17] unifies heterogeneous human behavior understanding tasks via language sequences and cycle training for each task in order. Nevertheless, these works are limited by the dilemma of complex training stages or task heads, which introduce extra costs. We integrate the output format of different tasks in skeleton sequence format and train the model just once.
@@ 3.1 Formulation
where $k$ represents the task index, while $i$ and $j$ indicate different example indexes. Thus, we can determine what task will be performed on the query sample by selecting the task-specific prompt from the corresponding training set.
@@
Figure 3: Overall framework of our Skeleton-in-Context. Specifically, we establish a skeleton bank by integrating training sets under different tasks, which contain a large amount of input-target pairs performing different tasks. Next, we randomly select a sample pair as the task-guided prompt (TGP) and a query input from the skeleton bank, undergo encoding and concatenating, respectively, and then input them into the transformer in parallel. In particular, during this process, the query input and task-unified prompt (TUP) are combined to form a new query. After iterating $n_{1}$ times, the TGP and query are aggregated through $\mathcal{A}\left(\cdot\right)$ and then input into the transformer for $n_{2}$ iterations. Lastly, the second half of the model output is used as our prediction.
@@ 3.2 Task Definitions
To measure multitasking capability, we establish a new in-context benchmark from common skeleton-based datasets, including H3.6M [24], AMASS [37], and 3DPW [56]. In this benchmark, we select three common tasks: motion prediction, pose estimation, and joint completion, and a new task: future pose estimation. Consequently, we conduct extensive experiments on a large and diverse dataset collection, which consists of 176K skeleton sequence samples.
Motion Prediction. Motion prediction [38]is to predict future motion based on historical motion. There is continuity between the input and output sequences, which constitute a complete motion. Thus, the input $I$ and target $T$ are complete skeleton sequences with shapes of $(F,J,3)$.
Pose Estimation. Pose estimation [18, 32, 34] is to lift 2D poses to 3D. The model takes the detected 2D skeleton sequence $I\in\mathbb{R}^{F\times J\times 2}$, which contain noise, as input, and estimates the accurate 3D skeleton sequence $T\in\mathbb{R}^{F\times J\times 3}$. Note that we expand the 2D input to 3D by filling the 3rd dimension with 00, in order to unify the data format with other tasks. Therefore, the input skeleton sequence becomes the same shape $(F,J,3)$ as the target sequence.
Future Pose Estimation. Based on prediction and estimation, we propose a new task: future pose estimation, which is to predict future 3D motion based on history 2D poses. As 2D images are easier to obtain than 3D data, this task unlocks a host of new real-world applications. Given a historical 2D skeleton sequence $I\in\mathbb{R}^{F\times J\times 3}$ (one dimension with 00), future pose estimation expects the model to output its future 3D skeleton sequence $T\in\mathbb{R}^{F\times J\times 3}$. To our knowledge, no work has ever proposed future pose estimation.
@@ 4 Skeleton-in-Context
4.1 Skeleton Sequence Prompt
Task-Guided Prompt. In in-context learning, a prompt is needed for the model to learn the underlying task pattern from context. As mentioned in Sec. 3.1, we propose the Task-Guided Prompt (TGP) $G^{k}_{i}=\left\{I^{k}_{i},T^{k}_{i}\right\}$, which comprises of an example skeleton pair performing a specific task $k$. In this way, our proposed model is able to perform a variety of tasks according to the provided prompts that could be seen or even unseen in the training set. It is worth noting that our model can perform unseen tasks and transfer to datasets that are different styles from the training dataset. It will be elaborated in Sec. 5.2.
Avoiding Over-Fitting. Previous works have made great progress in unifying different tasks via the MIM-style framework [5, 59, 16]. Compared with static objects in 2D and 3D, such as images and point clouds, the obstacle to in-context learning of skeleton sequences lies in the wondrous similarity of each frame, so directly using the random masking strategy [59, 5, 16] will cause the model to fall into over-fitting, leading it to only learn simple interpolation frames rather than capturing meaningful motion sequence representations during the training stage. As shown in Fig. 2, when we directly utilize the MPM-style [21, 44] framework to train our model, although it can perfectly reconstruct the masked frame during training, the experimental results reveal that it outputs a devastating result during inference. It is obvious that the output skeleton sequence gradually shrinks as the length increases because the next frame of the current frame is 0, and interpolating with 0 will lead to shrinking. Such a phenomenon underscores the challenges of learning contextual information about dynamic skeleton sequences.
@@ 4.1 Skeleton Sequence Prompt
siMLPe [41] WACV’23 19.0 30.0 35.3 46.8 52.4 36.7 74.6 60.0 59.0 67.4 63.2 64.4 68.4 66.4
EqMotion [67] CVPR’23 14.0 25.3 30.2 40.5 46.6 31.3 163.3 110.1 38.9 56.0 47.5 88.1 89.1 88.6
STCFormer [53] CVPR’23 17.4 28.0 31.5 43.0 40.4 33.9 55.6 45.6 33.6 50.6 42.1 62.2 71.7 67.0
GLA-GCN [72] ICCV’23 35.3 39.8 42.3 49.2 56.1 44.5 78.6 62.9 74.1 81.9 78.0 60.6 68.1 64.4
MotionBERT  [79] ICCV’23 14.3 22.8 26.7 36.4 41.7 28.4 55.8 45.7 31.3 46.7 39.0 75.5 96.8 86.2
@@
Task-Unified Prompt. With the Task-Guided Prompt, which exhibits unique capabilities specific to in-domain or out-of-domain tasks, mitigating the gap between different tasks presents a new challenge. Meanwhile, as discussed above, a straight extension of previous works will lead to a simple model that can only perform the interpolation task based on the given skeleton sequence. To tackle these challenges, we introduce the Task-Unified Prompt (TUP), a well-designed and task-agnostic module, which can adaptively learn to incorporate multiple tasks into the framework of in-context learning, i.e., a one-off end-to-end framework for all tasks without having to deploy task-specific heads or fine-tuning. We introduce two versions of TUP:
@@
We denote the SiC-based on Static-/Dynamic-TUP is referred to as Static-/Dynamic-SiC, respectively. The proposed TUP facilitates the modeling of distribution differences among tasks and datasets, so the model focuses more on learning unified representations shared across tasks.
@@ 4.2 Model Architecture
As shown in Fig. 3, we build a skeleton bank, a collection of publicly available datasets including H3.6M, AMASS and 3DPW, covering common skeleton analysis tasks, containing input-target pairs formed as sequences of equal length, where prompts are randomly selected. Given a prompt pair $G^{k}_{i}=\left\{I^{k}_{i},T^{k}_{i}\right\}$ and a query sample $Q^{k}_{j}$, we first map them to the C-dimensional feature space, and then splice $I$ and $T$, $Q$ and $U$ respectively. The former represents our proposed Task-Guided Prompt, and the latter is the objective sequence of our focused analysis. The process can be formulated as: